Find the soultion to the system of equations.
8x - 5y = 31
x - 4y = 14

Answer by unlockmath(1688)

Hello,
There's two to solve this. Let's do substitution method. We can get x alone in the second equation so it looks like:
x = 14+4y Now use this in the first equation to look like:
8(14+4y)-5y = 31 Combine like terms gives us:
112 + 27y=31 Subtract 112 from both sides gives us:
27y=-81 Divide both sides by 27 gives us:
y=-3 Now we can find x using either original equation.
8x - 5(-3) = 31
8x-15=31
x=2 Plug these into the original equation and it should come out good.
RJ Toftness
www.math-unlock.com